OpenStudy (anonymous):

describe a pattern in the sequence of numbers. what is the next number. 5, 0, -5, 10

OpenStudy (jamesj):

Did you mean 5, 0, -5, -10, ??? If so, this is an arithmetic progression, subtracting each time by 5. (or adding by -5 if you prefer.)

OpenStudy (anonymous):

is the fourth term supposed to be -10

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

if so the nth term can be found using\[a _{n}=5+(n-1)(-5)=10-5n\]

OpenStudy (anonymous):

is it -15

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

the so-called common difference (how you progress through the sequence) is -5; so the next term is -15

OpenStudy (anonymous):

so i add -5 to the numbers?

OpenStudy (anonymous):

yes